OpenStudy (fools101):

Same thing please!! y = 3x4 + 11x3 - 4x2 = x2(3x2 + 11x - 4)

OpenStudy (fools101):

please help!

OpenStudy (anonymous):

x^2= 0, x=0 3x^2+11x-4=(3x-1)(x+4)=0, x=-4, x=1/3